# ThinkTank: A Benchmark for Cognitive Intelligence in Pre-trained Language Models

### v1

@@ B.4.1 Scholastic Assessment Test analogy questions
Previous work has shown that fluid reasoning correlates with analogical reasoning GOSWAMI198673; snow1984topography; cattell1987intelligence. AI, ML, and NLP research has focused on analogical reasoning because this requires many componential abilities: syntactic parsing, semantic understanding, categorization, inductive reasoning, mathematical reasoning, and so on mil. Research on the cognitive alignment of PLMs has focused on performance on the 374 Scholastic Assessment Tests (SAT) analogy questions by turney2005measuring. Despite being broadly used in literature turney2005measuring; turney2010frequency; hendrickx2019semeval; webb2023emergent, our pilot experiments show that PLMs like GPT-3.5-Turbo, GPT-4, and Gemini perform nearly at ceiling on this test, while other open source models perform poorly on the same test. This hints that the set of questions in the test may be part of the GPT-X/ Gemini training or tuning data.

### v3

@@ B.4.1 Scholastic Assessment Test analogy questions
Previous work has shown that fluid reasoning correlates with analogical reasoning Goswami (1986); Snow et al. (1984); Cattell (1987). AI, ML, and NLP research has focused on analogical reasoning because this requires many componential abilities: syntactic parsing, semantic understanding, categorization, inductive reasoning, mathematical reasoning, and so on Pearson (2021). Research on the cognitive alignment of PLMs has focused on performance on the 374 Scholastic Assessment Tests (SAT) analogy questions by Turney (2005). Despite being broadly used in literature Turney (2005); Turney and Pantel (2010); Hendrickx et al. (2019); Webb et al. (2023), our pilot experiments show that PLMs like GPT-3.5-Turbo, GPT-4, and Gemini perform nearly at ceiling on this test, while other open source models perform poorly on the same test. This hints that the set of questions in the test may be part of the GPT-X/ Gemini training or tuning data.